# MV-MATH: Evaluating Multimodal Math Reasoning in Multi-Visual Contexts

### v5

@@ 1 Introduction
"The more you see, the more you know." — Leonardo da Vinci Large Language Models (LLMs) have demonstrated exceptional performance across a wide range of tasks including natural language understanding, question answering and code generation [31, 38, 49, 5, 29, 46]. Building on LLMs, MLLMs excel at various multimodal tasks by learning specialized features from different modalities, demonstrating a strong ability to understand different modalities, and have achieved remarkable results in various visual-language tasks such as image captioning, visual question answering, and visual reasoning [54, 44, 4, 8, 15, 42].
@@ 2.1 Benchmark for Mathematical Reasoning
All of the above datasets are restricted to single-visual contexts, relying on a single image input per problem. In contrast, our MV-MATH dataset emphasizes multi-visual mathematical reasoning by incorporating multiple images per problem. Addressing multi-image problems requires models to develop a more comprehensive understanding of relationships between mathematical visual elements, aligning more close with real-world mathematical scenarios.
@@ 2.2 Multimodal Foundation Models
The progress in visual-language alignment and rapid evolution of LLMs have endowed these models with visual capabilities, with notable achievements including CLIP [35] and BLIP [19]. CLIP aligns image-text pairs via contrastive learning, while BLIP combines contrastive and generative learning to align and generate image-text pairs. Inspired by these, MiniGPT-4 [55], LLaMA-Adapter [51], and LLaVA [24] extend text-only models into multimodal ones through vision-language alignment and instruction tuning. Subsequently, large multimodal models emerged, such as the closed-source GPT-4V [30], Gemini [37], Claude [2], GPT-4o [32] and open-source DeepSeek-VL [27], Qwen-VL [3], InternLMXComposer-VL [10], LLaVA-NeXT [23], and LLaVA-OneVision [17], which perform well on benchmarks like MMMU [48] and MathVista [28]. In mathematics, G-LLaVA [13] and Math-LLaVA [36] enhance performance on mathematical tasks using large instruction datasets Geo170K [13] and MathV360K [36]. MAVIS [53] improves visual perception with three-stage fine-tuning, while MultiMath [33] refines visual-language alignment through a four-stage training process.
@@ 3.1 Overview
Table 1: Key statistics of MV-MATH. The unit of question length is words.
@@ 3.2 Data Construction
Data labeling. Using the above samples, we conduct difficulty grading, subject categorization, and image relevance classification. For each problem, difficulty is categorized by analyzing the lengths of the question and analysis fields, assigning weights of $0.4$ and $0.6$ respectively. Based on the weighted lengths, MV-MATH is divided into three levels: easy, medium, and hard. For subject categorization and image relevance, we apply a majority voting approach, leveraging classification results from GPT-4o, Claude-3.5-Sonnet, and Qwen-VL-Max to determine each problem’s subject and its degree of image dependence. We then manually validate the assigned difficulty levels, subjects, and image relevance, with two annotators independently reviewing each problem and making adjustments to correct any misclassifications, ensuring the accuracy of the final classification.
Figure 3: Comparison of MV-MATH with other existing benchmarks. The size of each circle represents the richness of image distribution per question. The definition and calculation method of CV are provided in Section 3.3
@@ 4.1 Main Results
Challenging Nature of MV-MATH: The data presented in  Table 4 illustrates the inherent challenges of MV-MATH. Notably, the overall accuracy of the leading model, Claude-3.5-Sonnet, is only 33.9%, which falls significantly short of the human level accuracy of 76.5%. Among the open-source models, only LLaVA-OneVision-72B (LLaVA-OV-72B) achieves an accuracy above 20% on MV-MATH, while all others fall below this threshold. Notably, Math-LLaVA, a specialized mathematical MLLM, demonstrates limited performance on the MV-MATH dataset, with an accuracy of only 3.0%. Our analysis suggests that this underperformance stems from the distribution of its training data. Specifically, Math-LLaVA trained on the Math360K dataset do not effectively generalize to the MV-MATH dataset, underscoring a significant limitation in its generalization capabilities. A more detailed analysis can be found in the Appendix. This finding underscores a broader challenge for current mathematical MLLMs: the need to achieve robust performance on both in-distribution and out-of-distribution data. Future advancements should prioritize strengthening generalization capabilities over simply memorizing training data.
@@ 4.2 Analysis of Question type and Difficulty
Different Difficulty Levels.  Table 2 compares the performance of selected models across three difficulty levels. GPT-4o demonstrates significantly higher proficiency with a success rate of 40.3% in the "Easy" category. For the "Medium" category, Claude achieves the best performance at 37.5%. The performance disparity further decreases in the "Hard" category, where the leading Claude model reaches only 26.6% accuracy. As the difficulty level increases, the performance gap between models becomes smaller. Of particular note, Claude is the only model whose performance on medium difficulty exceeds that on low difficulty, whereas the performance of other models progressively declines as the difficulty level increases.
@@ 4.3 Impact of Image Relevance and Input Methods
Input Methods, Merge vs. Sequential: To investigate the impact of image input methods on model performance, we apply both merge and sequence input approaches on the MV-MATH dataset, with results presented in Figure 4(b). The findings consistently demonstrate that the sequence input method outperforms the merge input method across all tested models, indicating that preserving the positional and sequential information of images is crucial for effective multi-image mathematical reasoning. This superior performance of sequence input underscores the importance of structured visual information in enhancing model capabilities for interpreting and processing complex mathematical scenarios. An interesting observation is that, while the LLaVA-OV-SI-72B model underperforms compared to the LLaVA-OV-Chat-72B model on sequence input method, it outperforms the latter on merge input method. This shift may be attributed to the initial single-image training of LLaVA-OV-SI, which enhances its ability to retain distinct visual features within individual images when processed in a merged format.
@@ 5 Error Analysis and Model Insights
We conduct a detailed error analysis of the best-performing model, Claude-3.5-Sonnet. We randomly sample 100 errors, examine their distribution in detail, and provide corresponding qualitative examples. Based on this analysis, we classify the errors into five types, with the specific categories and their distribution shown in Figure 5.
@@ A.3 Division of Difficulty Levels
[…]
@@ B.4 Model Details
Step 2: So the area of this triangle is $3\times 4/2=6$.
@@ F.1 Data Collection
We have crawled a large number of multimodal math test questions from "Zujuan" totaling around 380k multimodal math questions. To format all questions for use, we process them by OCR engine like Mathpix interface. Due to inherent errors in the OCR engine, we introduce manual checks to ensure the accuracy of parsing results and to verify whether the questions belong to multimodal math problems. Specifically, we use the three-stage strategy outlined in  Section 3 to complete the final filtering of the data, ensuring not only the high quality of the questions but also the high quality of the images.
@@ H Case Study
In this section, we provide more detailed error examples of Claude-3.5-Sonnet. We classify Claude’s errors into five categories: Visual Perception Error, Reasoning Error, Calculation Error, Knowledge Error, and Reject Error. Detailed examples can be seen in Figures 21, 22, 23, 24, 25, 26 and 27.

### v6

@@ 1 Introduction
”The more you see, the more you know.” — Leonardo da Vinci Large Language Models (LLMs) have demonstrated exceptional performance across a wide range of tasks including natural language understanding, question answering and code generation [31, 38, 49, 5, 29, 46]. Building on LLMs, MLLMs excel at various multimodal tasks by learning specialized features from different modalities, demonstrating a strong ability to understand different modalities, and have achieved remarkable results in various visual-language tasks such as image captioning, visual question answering, and visual reasoning [54, 44, 4, 8, 15, 42].
@@ 2.1 Benchmark for Mathematical Reasoning
All of the above datasets are restricted to single-visual contexts, relying on a single image input per problem. In contrast, our MV-MATH dataset emphasizes multi-visual mathematical reasoning by incorporating multiple images per problem. Addressing multi-image problems requires models to develop a more comprehensive understanding of relationships between mathematical visual elements, aligning more closely with real-world mathematical scenarios.
@@ 2.2 Multimodal Foundation Models
The progress in visual-language alignment and rapid evolution of LLMs have endowed these models with visual capabilities, with notable achievements including CLIP [35] and BLIP [19]. CLIP aligns image-text pairs via contrastive learning, while BLIP combines contrastive and generative learning to align and generate image-text pairs. Inspired by these, MiniGPT-4 [55], LLaMA-Adapter [51], and LLaVA [24] extend text-only models into multimodal ones through vision-language alignment and instruction tuning. Subsequently, large multimodal models emerged, such as the closed-source GPT-4V [30], Gemini [37], Claude [2], GPT-4o [32] and open-source DeepSeek-VL [27], Qwen-VL [3], InternLMXComposer-VL [10], LLaVA-NeXT [23], and LLaVA-OneVision [17], which perform well on MMMU [48] and MathVista [28]. In mathematics, G-LLaVA [13] and Math-LLaVA [36] enhance performance on mathematical tasks using large instruction datasets Geo170K [13] and MathV360K [36]. MAVIS [53] improves visual perception with three-stage fine-tuning, while MultiMath [33] refines visual-language alignment through a four-stage training process.
@@ 3.1 Overview
Table 1: Key statistics of MV-MATH.
@@ 3.2 Data Construction
Data labeling. Using the above samples, we conduct difficulty grading, subject categorization, and image relevance classification. For each problem, difficulty is categorized by analyzing the lengths of the question and analysis, assigning weights of 0.40.40.4 and 0.60.60.6 respectively. Based on the weighted lengths, MV-MATH is divided into three levels: easy, medium, and hard. For subject categorization and image relevance, we apply a majority voting approach, leveraging classification results from GPT-4o, Claude-3.5-Sonnet, and Qwen-VL-Max to determine each problem’s subject and its degree of image dependence. We then manually validate the assigned difficulty levels, subjects, and image relevance, with two annotators independently reviewing each problem and making adjustments to correct misclassifications, ensuring the accuracy of the final result.
Figure 3: Comparison of MV-MATH with other existing benchmarks. The size of each circle represents the image distribution per question. The definition and calculation method of CV are provided in Section 3.3
@@ 4.1 Main Results
Challenging Nature of MV-MATH: The data presented in  Table 4 illustrates the inherent challenges of MV-MATH. Notably, the overall accuracy of the leading model, Claude-3.5-Sonnet, is only 33.9%, which falls significantly short of the human level accuracy of 76.5%. Among the open-source models, only LLaVA-OneVision-72B (LLaVA-OV-72B) achieves an accuracy above 20% on MV-MATH, while all others fall below this threshold. Notably, Math-LLaVA, a specialized mathematical MLLM, demonstrates limited performance on the MV-MATH dataset, with an accuracy of only 3.0%. Our analysis suggests that this underperformance stems from the distribution of its training data. Specifically, Math-LLaVA trained on the Math360K dataset do not effectively generalize to the MV-MATH dataset, underscoring a significant limitation in its generalization capabilities. This finding underscores a challenge for current mathematical MLLMs: the need to achieve robust performance on both in-distribution and out-of-distribution data. Future advancements should prioritize strengthening generalization capabilities over memorizing training data.
@@ 4.2 Analysis of Question type and Difficulty
Different Difficulty Levels.  Table 2 compares the performance of selected models across three difficulty levels. GPT-4o demonstrates significantly higher proficiency with a success rate of 40.3% in the ”Easy” category. For the ”Medium” category, Claude achieves the best performance at 37.5%. The performance disparity further decreases in the ”Hard” category, where the leading Claude model reaches only 26.6% accuracy. As the difficulty level increases, the performance gap between models becomes smaller. Of particular note, Claude is the only model whose performance on medium difficulty exceeds that on low difficulty, whereas the performance of other models progressively declines as the difficulty level increases.
@@ 4.3 Impact of Image Relevance and Input Methods
Input Methods, Merge vs. Sequential: To investigate the impact of image input methods on model performance, we apply both merge and sequence input approaches on MV-MATH , with results presented in Figure 4(b). The findings demonstrate that the sequence input method outperforms the merge input method across all tested models, indicating that preserving the positional and sequential information of images is crucial for effective multi-image mathematical reasoning. This superior performance of sequence input underscores the importance of structured visual information in enhancing model capabilities for interpreting and processing complex mathematical scenarios. An interesting observation is that, while the LLaVA-OV-SI-72B model underperforms compared to the LLaVA-OV-Chat-72B model on sequence input method, it outperforms the latter on merge input method. This shift may be attributed to the initial single-image training of LLaVA-OV-SI, which enhances its ability to retain distinct visual features within individual images when processed in a merged format.
@@ 5 Error Analysis and Model Insights
We conduct a detailed error analysis of Claude-3.5-Sonnet. We randomly sample 100 errors, examine their distribution in detail, and provide corresponding qualitative examples. Based on this analysis, we classify the errors into five types, with the specific categories and their distribution shown in Figure 5.
@@ 6 Conclusion
Acknowledgment
This work has been supported by the National Natural Science Foundation of China (NSFC) Grant 62436009.
@@ A.3 Division of Difficulty Levels
Step 1: A weighted average is calculated using the lengths of the question and analysis fields, assigning a weight of 0.4 to the question length and 0.6 to the analysis length. This weighting reflects the assumption that the solution process (captured in the analysis field) is more indicative of a question’s difficulty. The weighted length distribution is shown in Figure 8. Using the K-means clustering algorithm, we cluster the weighted lengths into three difficulty levels: easy (01500–1500 – 150), medium (150500150–500150 – 500), and hard (>500>500> 500).
@@ B.4 Model Details
Step 2: So the area of this triangle is 3×4/2=63$\times$ 4/2=63 × 4 / 2 = 6.
@@ F.1 Data Collection
We have crawled a large number of multimodal math test questions totaling around 380k multimodal math questions. To format all questions for use, we process them by OCR engine like Mathpix interface. Due to inherent errors in the OCR engine, we introduce manual checks to ensure the accuracy of parsing results and to verify whether the questions belong to multimodal math problems. Specifically, we use the three-stage strategy outlined in  Section 3 to complete the final filtering of the data, ensuring not only the high quality of the questions but also the high quality of the images.
@@ H Case Study
In this section, we provide more detailed error examples of Claude-3.5-Sonnet. We classify Claude’s errors into five categories: Visual Perception Error, Reasoning Error, Calculation Error, Knowledge Error, and Reject Error. Detailed examples can be seen in Figures 21, 22, 23, 24, 25, 26, 27, 28, 29 and 30.
@@
Figure 22: An example of visual perception error.
Figure 23: An example of reasoning error.